OpenStudy (anonymous):

The radius of a sphere is increasing at a rate of 3 mm/s. How fast is the volume increasing when the diameter is 80 mm?

OpenStudy (anonymous):

\[Volume=\frac{4}{3}\pi r^3\] In terms of the diameter (\(D=2r\)): \[V=\frac{4}{3}\pi\left(\frac{D}{2}\right)^3\\ V=\frac{1}{6}\pi D^3\] Implicit differentiation: \[\frac{dV}{dt}=\frac{1}{2}\pi D^2\frac{dD}{dt}\] The radius is increasing at 3 mm/s, so \[\frac{dr}{dt}=3~\Rightarrow~\frac{dD}{dt}=6\] Now plug in \(\dfrac{dD}{dt}\) and the given \(D=80~mm\): \[\frac{dV}{dt}=\frac{1}{2}\pi (80~mm)^2\left(6\frac{mm}{s}\right)\]